40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Wed Jan 30 02:20:38 2013, tara, Notes, Drawings, layout for the new setup 
    Reply  Fri Feb 1 03:14:45 2013, tara, Notes, Drawings, layout for the new setup ctn_layout_2laser.pdf
       Reply  Mon Feb 18 03:16:44 2013, tara, Notes, Drawings, layout for the new setup ctn_layout_2laser_v2.png
          Reply  Fri Jun 21 21:02:40 2013, tara, Notes, Drawings, layout for the new setup 
             Reply  Mon Jun 24 21:27:02 2013, tara, Notes, RefCav, setup for 2nd refcav 
Message ID: 1098     Entry time: Fri Feb 1 03:14:45 2013     In reply to: 1096     Reply to this: 1100
Author: tara 
Type: Notes 
Category: Drawings 
Subject: layout for the new setup 

Quote:

I'm making a new layout for the 2 laser setup. It is in progress. I just want to make a note about some features I want to have in the setup.

  • Move EAOM (for intensity stabilization,ISS) some where down stream, close to the chamber. Since the monitor is behind the cavity.
  • Pick up for optical fiber: This should be somewhere behind 1)EAOM, so that power can be stabilized, 2) BB EOM, so that the laser frequency is stabilized at higher frequency as well. 
  • Some beam splitter/ flip mirrors to bring laser to both cavities. so that I can scan two cavities with 1 laser at the same time. This should be handy when we need to find the beat, since the beat frequency is not well determined (by AOM) anymore.
  • Some space for an independent frequency stabilized laser. This is for the future, when we want to lock a laser to the 8" cavity and distribute it to other labs.  From the current setup, I think we might have to borrow the laser from the 2-laser setup, after the PMC, then lock it to the 8" cavity. Installing 3 sets of laser, PMC, refcav on the table is not easy. With this idea, we can lock the laser to a low noise cavity(8"), and just change the cable, when we want to switch to do the coating measurement.

 The tentative plan is shown below. I also move the BB eom for locking frequency(phase) next to the laser.

  • The pick up for fiber distribution can be at the PBS just after the PMC where we dump the power. However the laser is not intensity stabilized yet, but I'll worry about it later.
  • I add pick up for RFAM monitor. It is the unused beam after the EAOM. The power will fluctuate when ISS is active, but should be fine for RFAM monitoring purpose.

I'm doing mode matching calculation to see if the lenses can be placed in the available area. Otherwise I need to fold the beam a few more times.

I also started removing the acoustic box, some optics on the table so that I can try to position the real optics to see if they can fit on the table.

 

ctn_layout_2laser.pdf

ELOG V3.1.3-